an angle equals 3x+12. i need to find the complement and supplement
Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

90-3x+12=
3x=102
x=34 degrees
90-34=56 degrees
180-34=146 degrees

Answer by solver91311(24713)



If two angles are complementary, then the sum of their measures is 90 degrees or radians. So the complement of an angle that measures is degrees or radians. The sum of the measures of supplementary angles is 180 degrees or radians, otherwise the expressions are the same.


John